OpenStudy (jessiegonzales):

MEDAL!! PLEASE CHECK MY ANSWER 8. The lateral area of a cone is 574pi cm^2. The radius is 29 cm. What is the slant height to the nearest tenth of a centimeter? A. 9.9cm B. 6.3cm C. 19.8cm<---- D. 12.6cm

OpenStudy (anonymous):

You are correct :)

OpenStudy (anonymous):

\[574 \pi = \pi r l = \pi (29) l\] \[l \approx 19.7931\]

OpenStudy (jessiegonzales):

thanks :)